OpenStudy (anonymous):

Solve for (x,y) {X+2y=3 {2x+3y=3

OpenStudy (akashdeepdeb):

As there are 2 equations! They will have 1 solution ONLY So let us multiply the first equaition with 2 to get x co-eff. common okay? It'll become = 2x + 4y = 6 Now subtract this equation by 2x + 3y = 3 So 2x + 4y - 2x - 3y = 6-3 so y would be? And then substitute y's value in any of the above equations to get the x value and then you will get the point of there intersection Understood? :)

OpenStudy (anonymous):

Why do you multiply by 2 on the first step

OpenStudy (akashdeepdeb):

There ARE actually many methods to do this! You could have either found x value in terms of y and then you could have substituted that value of x in the second equation! But in the multiplication method! We should get the co-efficient of the variable x OR y [In this case 2 for x ] Common so that when we subtract OR add the 2 equations IT GET'S CUT!!

OpenStudy (akashdeepdeb):

For example in our case we could have also multiplied the first equation with 3 so that the y co-eff. in the first equation would become 6! But then we would have had to multiply the 2nd equation also with 2 so that the y co-eff would become 6! So that when we subtractted the 2 equations y would get CUT!! :-)

OpenStudy (akashdeepdeb):

Getting the concept?

OpenStudy (anonymous):

Yeah I got it. Thank you :)

OpenStudy (anonymous):

Want to hhelp me with one more

OpenStudy (akashdeepdeb):

Do you need help?

OpenStudy (anonymous):

Yeah this other one is confusing

OpenStudy (akashdeepdeb):

Kudos me? :)

OpenStudy (anonymous):

{2x-3(y+1)=8 {3(x+2)+5y=-6

OpenStudy (akashdeepdeb):

Okay so even in this case what you want to do is you've got to open the brackets for BOTH the equations using distributive property Al right? Tell me what you get then!

OpenStudy (anonymous):

I dont knowhow

OpenStudy (akashdeepdeb):

So it goes like this Say you have -ax(bx - c) Now how will you open the brackets here?

OpenStudy (akashdeepdeb):

TRY TRY!! :D

OpenStudy (anonymous):

{2x-3y+3=8 {3x+6+5y=-6

OpenStudy (akashdeepdeb):

Second one is CORRECT but the first one should be 2x - 3y - 3!! Getting why? :)

OpenStudy (anonymous):

Yeah I forgot about the -

OpenStudy (akashdeepdeb):

Okay so now 2x - 3y - 3 = 8 right? Now bring the -3 to the RHS it becomes 2x - 3y = 11 Do the same with the second equation now!

OpenStudy (anonymous):

3x+5y=0

OpenStudy (anonymous):

Then what?

OpenStudy (akashdeepdeb):

No, the second one would become NEGATIVE when shifted there right?

OpenStudy (anonymous):

Oh -12

OpenStudy (anonymous):

Okay So do I subtract the two equations?

OpenStudy (anonymous):

Hello?

OpenStudy (akashdeepdeb):

Hold on

OpenStudy (akashdeepdeb):

Okay so the 2 equations are? 2x - 3y = 11 3x+5y=-12 Now see for x and y! Do you see the co-eff of x in both equations same? Or do you see the co-effecient of y in both equations same? YES/NO? :)

OpenStudy (anonymous):

No?

OpenStudy (anonymous):

Am I supposed to?

OpenStudy (akashdeepdeb):

You are right IT IS NO!! So that is why we MULTIPLY [*HURRAY*]

OpenStudy (akashdeepdeb):

Do you understand what co-eff. means?

OpenStudy (anonymous):

No

OpenStudy (akashdeepdeb):

Okay so let us say we have ax Now the co-efficient is like the constant attached to the variable Okay? So in here the( co-efficient of x) = a Got that?

OpenStudy (anonymous):

I think so

OpenStudy (akashdeepdeb):

\[1x/2 , 4x/5 , c^2x\] Find the co-efficients of these! TRY! I'll help you

OpenStudy (akashdeepdeb):

There are 3 of them okay?

OpenStudy (anonymous):

Are they part of the problem

OpenStudy (akashdeepdeb):

Not our problem! I am trying to help you understand a very key part to solve your question!

OpenStudy (akashdeepdeb):

Do you want to do the 3 I gave you or not? Because I CAN tell you an easier way to solve your question! But what I am showing you now is really very helpful! YOUR CHOICE!

OpenStudy (anonymous):

So it's like anything thats paired together that aren't the variables

OpenStudy (akashdeepdeb):

YES!!! :D

OpenStudy (anonymous):

K :)

OpenStudy (akashdeepdeb):

let us get back to our question now

OpenStudy (akashdeepdeb):

2x - 3y = 11 3x+5y=-12 Tell me the co-efficient of x in both the equations Tell me the co-efficient of y in both the equations [INCLUDE THE NEGATIVE SIGNS IF IT IS THERE]

OpenStudy (anonymous):

2 and 3 3 and 5

OpenStudy (anonymous):

Er...-3

OpenStudy (akashdeepdeb):

Slightly wrong! The co-eff. of y in the first equation is -3!! Get it??

OpenStudy (akashdeepdeb):

YES!!! :D

OpenStudy (anonymous):

K so I multiply those

OpenStudy (akashdeepdeb):

Now we have to make one of the variables' co-eff. SAME in both equations! Getting me?

OpenStudy (anonymous):

Yeah I got you but how do I do that

OpenStudy (akashdeepdeb):

Good now let us say we want the co-eff. of x same in both the equations! okay?

OpenStudy (anonymous):

K

OpenStudy (akashdeepdeb):

So the co-eff of x are 2 3 right?

OpenStudy (akashdeepdeb):

Now we want them to be the same! So we have to multiply them with some numbers that will fetch us similar numbers! Do you know how to find LCM?

OpenStudy (anonymous):

I think so. Can we pick up the pace a bit. I have school in 7 hours and still need to get to bed:/

OpenStudy (akashdeepdeb):

Sure! Now see if we have 2 and 3 We kinda need to find a number to multiply it with to get something common! And that common number for the first equation is 3 And for the 2nd equation is 2! So it is kinda like finding the LCM! And when we multiply the equations accordingly we'd get 6x in both! Okay?

OpenStudy (anonymous):

K

OpenStudy (akashdeepdeb):

Like this 3*(2x - 3y = 11) = 6x - 9y = 33 2*(3x+5y=-12) = 6x + 10y = -24 Now we do this because if we have something common in both equations then we can cut them off when we subtract them!

OpenStudy (akashdeepdeb):

So coming back to your question! {X+2y=3 --> multiply with 2 to get co-eff of x 2 in both cases! {2x+3y=3 --> multiply with nothing because we already have the co=eff for x as 2! And then subtract the 2 equations Liek 2x + 4y - 2x - 3y = 6-3 (CHOP THE 2x off) You get y=3 And then put the y value in the first equation [x + 2y = 3] x + 2(3) = 3 x = -3 So the point will be (x,y) = (-3,3) tHE ANSWER TO YOUR QUESTION!! :) Getting me?? :D

OpenStudy (akashdeepdeb):

@245424 ?

OpenStudy (anonymous):

I got confusedt

OpenStudy (akashdeepdeb):

Then do this easier thing! X+2y=3 -> x = 3 -2y {2x+3y=3 Now substitute x's value in the 2nd equation as 2(3-2y) + 3y = 3!! And now open the brackets and solve! Getting that??

OpenStudy (anonymous):

I think we messed up. My teacher gave us the answers. He got (-1,3)

OpenStudy (akashdeepdeb):

Either you wrote the questionwrong! Or your teacher was wrong! Tell your teacher [TRY NOT TO BE DISRESPECTFUL] THAT HE WAS WRONG!

OpenStudy (akashdeepdeb):

Tell him IT IS (-3,3)

OpenStudy (anonymous):

Oh well. He said as long as we tried he would give us credit. So I'll keep this work and just deal with it. Thank you vermouth for helping me

OpenStudy (anonymous):

*very much

OpenStudy (akashdeepdeb):

Best response maybe?? :!

OpenStudy (akashdeepdeb):

Only RESPONSE maybe!! :P

OpenStudy (anonymous):

How do I do that

OpenStudy (akashdeepdeb):

Just press best response beside my answers!

OpenStudy (akashdeepdeb):

Have fun in Math class! Tell your sir to reCHECK!

OpenStudy (anonymous):

K thanks :)

OpenStudy (akashdeepdeb):

You're most wecome!